Jeff is two years younger than Carrie. 12 years ago, Carrie was twice as old as Jeff. How old is Jeff now?
Thank you!:o)

Answer by checkley77(12844)

J=C-2
J-12=2(C-12)
(C-2)-12=2C-24
C-2-12=2C-24
C-2C=-24+2+12
-C=-10
C=10 YEARS OLD NOW.
J=10-2=8 YEARS OLD NOW.
PROOF:
8-12=2(10-12)
-4=2*-2
-4=-4